# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

### v1

@@ 2.4.1 Employed with $np$ firm
Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in an $np$ firm. The worker chooses consumption $c$ and future savings $a^{\prime}$ to maximize their lifetime discounted utility. Their income constitutes of period wage $\omega_{np}$ net of payroll tax $\tau$, current savings $(1+r)a$, and dividends $d$. The expectation of the worker’s future value is taken over the idiosyncratic productivity shock $s$ and the race-specific probability of extreme wealth shock $\epsilon_{R}$.
@@
The future value of an employed worker for each $s^{\prime}$ and after realizing $\epsilon_{R}$ can expand to Equation 3.
@@
If the worker receives a job destruction shock, specific to $np$ firms, one becomes unemployed with probability $\lambda_{np}$. Otherwise, one may stay employed with probability $1-\lambda_{np}$. If unemployed, the worker can receive unemployment benefits and realize a value of $U^{I}_{np}$.
@@ 2.4.2 Employed with $p$ firm
Prejudiced firms $p$ hire only white workers. The rest of the problem is similar to the previous case and is given by Equation 4.
@@ 2.4.3 Unemployed and eligible workers
Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or an $np$ firm. In addition, unemployed white workers can receive job offers from both $p$ and $np$ firms, while black workers can get matched only with $np$ firms.
Black worker from a $np$ firm
Equation 6 describes the value of an unemployed black worker eligible for benefits. The worker receives unemployment insurance payout, $b_{np}(bl,s,a)$, net of labor income tax rate $\tau$, previous savings and lump-sum dividend transfer $d$. The worker encounters the extreme wealth shock with probability $\epsilon_{bl}$.
@@
The unemployed worker finds a job with probability $f(\theta_{np})$. If the worker doesn’t get matched, they either continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $1-P_{e}$.
White worker from a $np$ firm
Unemployed white workers are recruited by both $p$ and $np$ firms. Their value function is described by Equation 8. Like a black worker, a white worker has cash-on-hand from post-tax unemployment insurance payout, previous savings, and a lump-sum dividend transfer. They also encounter an extreme wealth shock with probability $\epsilon_{wh}$.
@@
Equation 9 describes the future value of an unemployed white worker with idiosyncratic productivity $s^{\prime}$ and asset $a^{\prime}$. They can receive a job offer from a $p$ firm with probability $(1-f(\theta_{np}))f(\theta_{p})$, an offer from an $np$ firm with probability $f(\theta_{np})(1-f(\theta_{p}))$, and receive both offers together with probability $f(\theta_{np})f(\theta_{p})$. The white worker will remain unemployed if they don’t receive any offers (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$) and continue to receive unemployment benefits with probability $P_{e}$. If the worker receives both $p$ and $np$ offers, they will choose the job with higher lifetime utility.
$\hat{U^{I}}_{np}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (9)
$+\underbrace{f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth p and np jobs}}$
$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))}_{\text{not matching with any job}}\Big{[}\underbrace{P_{e}U^{I}_{np}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing UI}}+\underbrace{(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of losing UI}}\Big{]}$
White worker from a $p$ firm
@@
After realizing $\epsilon_{wh}$ and for each $s^{\prime}$, the worker’s future value is described in Equation 11. It is nearly identical to Equation 9 except for the value of remaining unemployed from a $p$ firm rather than a $np$ firm.
$\hat{U^{I}}_{p}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (11)
$+\underbrace{f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth p and np jobs}}$
$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))}_{\text{not matching with any job}}\Big{[}\underbrace{P_{e}U^{I}_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing UI}}+\underbrace{(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of losing UI}}\Big{]}$
@@ 2.4.4 Unemployed and ineligible workers
Ineligible workers have lost their unemployment insurance and are still searching for jobs. They no longer receive unemployment benefits unless they successfully find a job and transition out of employment again. Thus, their income comes only from their previous savings and lump-sum dividends and does not depend on their prior employment situation.
@@
Equation 12 describes the problem of a black worker searching for jobs without unemployment insurance.
@@
And Equation 13 expands the future value of the black worker.
@@
And Equation 15 expands the future value of the white worker in this scenario.
$\hat{U^{N}}(wh,s^{\prime},a^{\prime})$ $=\underbrace{(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})}_{\text{value of matching with a p sector job only}}+\underbrace{f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})}_{\text{value ofmatching with a np sector job only}}$ (15)
$+\underbrace{f(\theta_{p})f(\theta_{np})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}}_{\text{value of matching withboth np and p sector jobs}}$
$+\underbrace{(1-f(\theta_{p}))(1-f(\theta_{np}))U^{N}(wh,s^{\prime},a^{\prime})}_{\text{value of continuing unemployed}}$
Thus, the problem faced by an ineligible worker is analogous to that faced by an eligible worker, except that they don’t receive unemployment benefits.
@@ 2.5 Firms
There are a large number of $p$ and $np$ firms that post vacancies after paying the vacancy posting costs. The firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
@@ 2.5.1 Vacant $np$ firm
$V_{np}$ $=-\kappa_{np}+\Big{(}\frac{q(\theta_{np})}{1+r}\Big{)}\int_{a}\Big{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\underbrace{J_{np}(bl,s^{\prime},a^{\prime})\frac{\phi_{u}(bl,s,a)}{u}}_{\text{value of matching with atype (s,a) black worker}}\Big{]}$ (16)
$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\underbrace{\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}}}_{\text{np provides higher value than a p firm}}\Big{(}\underbrace{J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}}_{\text{value of matching with (s,a) white worker}}\Big{)}$
$+\underbrace{\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}}_{\text{p provides higher value than a np firm}}\Big{(}\underbrace{J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}(1-f(\theta_{p}))}_{\text{value of matching (s,a) worker without competing p offer}}\Big{]}\Big{\}}da,$
The firm’s expected value of matching with a worker depends on the probability of matching with a specific worker of type $(R,s,a)$ and the operating value, $J_{np}(R,s,a)$. A $np$ firm can match with either a black or a white worker from the current unemployment pool. $\phi_{u}(bl,s,a)$ is the mass of unemployed black workers with productivity $s$ and asset $a$, while $\phi_{u}(wh,s,a)$ is the corresponding mass of white workers. Thus, $\phi_{u}(bl,s,a)/u$ and $\phi_{u}(wh,s,a)/u$ are the densities of the unemployed black and white workers over $s$ and $a$. These reflect the probabilities that a vacant $np$ firm gets matched with an unemployed worker of type $(R,s,a)$. Since the white unemployed workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ match and begin producing only if it gives them a higher value or if the worker did not receive a $p$ firm offer. The last two rows of Equation 16 reflect the cases that a white worker will produce with the $np$ firm.
@@ 2.5.2 Vacant $p$ firm
Vacant $p$ firms pay a posting cost of $\kappa_{p}$ and restrict their search to unemployed white workers. The probability of matching with a white unemployed worker is $q(\theta_{p})$. Equation 17 describe the value of maintaining a $p$ vacancy, $V_{p}$.
$V_{p}$ $=-\kappa_{p}+\Big{(}\frac{q(\theta_{p})}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\int_{a}\Big{[}\underbrace{\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})>W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}}_{\text{value of matching with (s,a) white worker, $p$ offers higher value than a np firm}}$ (17)
$+\underbrace{\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})\leq W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}(1-f(\theta_{np}))}_{\text{value of matching with (s,a) white worker without competing np offer}}\Big{]}da$
Since the vacant $p$ firm searches only among white workers, the probability of matching with an unemployed white worker having productivity $s$ and asset $a$ is given by $\phi_{u}(wh,s,a)/u_{wh}$. Like the case of $np$ firms, the white worker accepts the $p$ match only if the $p$ offer is more favorable or if they don’t receive a competing $np$ offer.
@@ 2.5.3 Producing $np$ firm
The value of a producing $np$ firm is described by Equation 18. $j(R,s,a)$ describes the period profit. The worker’s productivity $s$ enters the production function. The matched firm produces $sf(k)$ units of output, pays the rental and the depreciation cost of capital, $(r+\delta)k$, and the wage cost, $\omega_{np}(R,s,a)$.
$J_{np}(R,s,a)$ $=\max_{k}\Big{\{}j(R,s,a)+\Big{(}\frac{1-\lambda^{np}}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{R})J_{np}(R,s^{\prime},a^{\prime})+\epsilon_{R}J_{np}(R,s^{\prime},0)\big{]}\Big{\}}$ (18)
@@
The firm discounts its future value using the interest rate adjusted for the job destruction shock $\lambda_{np}$. Since the worker may experience an extreme wealth shock with probability $\epsilon_{R}$, it changes the wage bargaining position of the worker to the firm. The future value of the firm incorporates this probability.
@@ 2.6 Producing $p$ firm
Similar to a producing $np$ firm, Equation 19 describes a producing $p$ firm. The firm hires capital and discounts future profit accounting for the job destruction shock $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.

### v2

@@ 2.4.1 Employed with np firm
Equation 2 describes an employed worker of race $R$ with productivity $s$ and asset holdings $a$, working in a $np$ firm. The worker chooses consumption $c$ and future savings $a^{\prime}$ to maximize their discounted lifetime utility. Their income constitutes of period wage $\omega_{np}$ net of payroll tax $\tau$, current savings $(1+r)a$, and dividends $d$. The expectation of the worker’s future value is taken over the idiosyncratic productivity shock $s$ and the race-specific probability of extreme wealth shock $\epsilon_{R}$.
@@
The future value of an employed worker for each $s^{\prime}$ and after realizing $\epsilon_{R}$ can expand to Equation 3. If the worker receives a job destruction shock, specific to $np$ firms, one becomes unemployed with probability $\lambda_{np}$. Otherwise, one may stay employed with probability $1-\lambda_{np}$. If unemployed, the worker can receive unemployment benefits and realize a value of $U^{I}_{np}$.
@@ 2.4.2 Employed with p firm
Equation 4 describes the case of a worker employed by a $p$ firm. It is analogous to the previous case, and we only keep track of white workers since prejudiced firms do not hire black workers.
@@ 2.4.3 Unemployed and eligible workers
Since unemployment insurance is proportional to the counterfactual wage, the value obtained by an unemployed and eligible worker depends on whether the past employment was in a $p$ or a $np$ firm. In addition, unemployed white workers can receive job offers from both $p$ and $np$ firms, while black workers can get matched only with $np$ firms.
Unemployed black worker from $np$ firm
Equation 6 describes the value of an unemployed black worker eligible for benefits. The worker encounters the extreme wealth shock with probability $\epsilon_{bl}$. The worker receives unemployment insurance payout, $b_{np}(bl,s,a)$, net of labor income tax rate $\tau$, previous savings, and lump-sum dividend transfer $d$.
@@
The unemployed worker finds a job with probability $f(\theta_{np})$, which moves them to a future value of $W_{np}(bl,s^{\prime},a^{\prime})$. Otherwise, they either continue to receive unemployment benefits with probability $P_{e}$ or lose their benefit with probability $(1-P_{e})$. If one loses the benefit, their value function becomes $U^{N}(bl,s^{\prime},a^{\prime})$.
Unemployed white worker from $np$ firm
Unemployed white workers are recruited by both $p$ and $np$ firms. Equation 8 describes their value function. Like a black worker, a white worker encounters an extreme wealth shock with probability $\epsilon_{wh}$. They also have a budget constraint from post-tax unemployment insurance payout, previous savings, and a lump-sum dividend transfer.
@@
Equation 9 describes the future value of an unemployed white worker with idiosyncratic productivity $s^{\prime}$ and asset $a^{\prime}$. With probability $(1-f(\theta_{np}))f(\theta_{p})$, the worker receives an only job offer from a $p$ firm. With probability $f(\theta_{np})(1-f(\theta_{p}))$, they receive an only offer from a $np$ firm. With probability $f(\theta_{np})f(\theta_{p})$, they receive offers from both $np$ and $p$ firms simultaneously. If the worker receives both $p$ and $np$ offers, they will choose the job with a higher lifetime value. The white worker will remain unemployed if they don’t receive any offers (with probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$). In this case, one may continue to receive unemployment benefits with probability $P_{e}$. Otherwise, one moves to a state of receiving no unemployment benefit with probability $(1-P_{e})$.
$\hat{U^{I}}_{np}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (9)
$+f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}$
$+(1-f(\theta_{p}))(1-f(\theta_{np}))\Big{[}P_{e}U^{I}_{np}(wh,s^{\prime},a^{\prime})+(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})\Big{]}$
Unemployed white worker from $p$ firm
@@
After realizing $\epsilon_{wh}$ and for each $s^{\prime}$, the unemployed worker’s future value is described in Equation 11. It is analogous to Equation 9.
$\hat{U^{I}}_{p}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (11)
$+f(\theta_{np})f(\theta_{p})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}$
$+(1-f(\theta_{p}))(1-f(\theta_{np}))\Big{[}P_{e}U^{I}_{p}(wh,s^{\prime},a^{\prime})+(1-P_{e})U^{N}(wh,s^{\prime},a^{\prime})\Big{]}$
@@ 2.4.4 Unemployed and ineligible workers
Ineligible workers have lost their unemployment insurance and are still searching for jobs. They no longer receive unemployment benefits unless they successfully find a job and transition out of employment again. Therefore, their consumption relies solely on previous savings and lump-sum dividends, independent of their prior employment status.
@@
Equation 12 describes the problem of an unemployed black worker searching for jobs without unemployment insurance.
@@
And Equation 13 expands the future value of the black worker. One has a probability of $f(\theta_{np})$ of finding a job - only with a $np$ firm; otherwise, they remain unemployed and ineligible.
@@
And Equation 15 expands the future value of the white worker in this scenario. It is nearly identical to Equation 11. With a probability $(1-f(\theta_{p}))(1-f(\theta_{np}))$, the unemployed worker does not find any job and remains unemployed and ineligible.
$\hat{U^{N}}(wh,s^{\prime},a^{\prime})$ $=(1-f(\theta_{np}))f(\theta_{p})W_{p}(wh,s^{\prime},a^{\prime})+f(\theta_{np})(1-f(\theta_{p}))W_{np}(wh,s^{\prime},a^{\prime})$ (15)
$+f(\theta_{p})f(\theta_{np})\max\{W_{p}(wh,s^{\prime},a^{\prime}),W_{np}(wh,s^{\prime},a^{\prime})\}$
$+(1-f(\theta_{p}))(1-f(\theta_{np}))U^{N}(wh,s^{\prime},a^{\prime})$
@@ 2.5 Firms
A large number of $p$ and $np$ firms post vacancies after paying the vacancy posting costs. These firms are risk-neutral and discount their future profits using the equilibrium real interest rate.
@@ 2.5.1 Vacant np firm
$V_{np}$ $=-\kappa_{np}+\Big{(}\frac{q(\theta_{np})}{1+r}\Big{)}\int_{a}\Big{\{}\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}J_{np}(bl,s^{\prime},a^{\prime})\frac{\phi_{u}(bl,s,a)}{u}\Big{]}$ (16)
$+\sum_{s^{\prime}}\pi_{ss^{\prime}}\Big{[}\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}$
$+\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{np}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u}\Big{)}(1-f(\theta_{p}))\Big{]}\Big{\}}da,$
The firm’s expected value of matching with a worker depends on the probability of matching with a specific worker of type $(R,s,a)$ and the operating value, $J_{np}(R,s,a)$. A $np$ firm can match with either a black or a white worker from the current unemployment pool. $\phi_{u}(bl,s,a)$ is the mass of unemployed black workers with productivity $s$ and asset $a$, while $\phi_{u}(wh,s,a)$ is the corresponding mass of white workers. Thus, $\phi_{u}(bl,s,a)/u$ and $\phi_{u}(wh,s,a)/u$ are the densities of the unemployed black and white workers over $s$ and $a$. These reflect the probabilities that a vacant $np$ firm gets matched with an unemployed worker of type $(R,s,a)$. Since unemployed white workers can potentially receive a competing offer from a $p$ firm, the white worker will accept the $np$ job and begin producing only if the $np$ job provides the worker with a higher value, described by $(\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})\geq W_{p}(wh,s^{\prime},a^{\prime})\}})$, or if the worker did not receive a $p$ firm offer, described by $(\mathbbm{1}_{\{W_{np}(wh,s^{\prime},a^{\prime})<W_{p}(wh,s^{\prime},a^{\prime})\}}(1-f(\theta_{p})))$.
@@ 2.5.2 Vacant p firm
Vacant $p$ firms pay a posting cost of $\kappa_{p}$ and restrict their search to unemployed white workers. The probability of matching with an unemployed white worker is $q(\theta_{p})$. Equation 17 describe the value of maintaining a $p$ vacancy, $V_{p}$.
$V_{p}$ $=-\kappa_{p}+\Big{(}\frac{q(\theta_{p})}{1+r}\Big{)}\sum_{s^{\prime}}\pi_{ss^{\prime}}\int_{a}\Big{[}\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})>W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}$ (17)
$+\mathbbm{1}_{\{W_{p}(wh,s^{\prime},a^{\prime})\leq W_{np}(wh,s^{\prime},a^{\prime})\}}\Big{(}J_{p}(wh,s^{\prime},a^{\prime})\frac{\phi_{u}(wh,s,a)}{u_{wh}}\Big{)}(1-f(\theta_{np}))\Big{]}da$
Since the vacant $p$ firm only searches among white workers, the probability of matching with an unemployed white worker who has productivity $s$ and assets $a$ is represented by $\phi_{u}(wh,s,a)/u_{wh}$. Similar to the case of $np$ firms, a worker will accept the $p$ job only if the offer is more favorable than other options or if they do not receive a competing offer from an $np$ firm.
@@ 2.5.3 Producing np firm
The value of a producing $np$ firm is described by Equation 18. The firm chooses capital $k$ to maximize its lifetime value producing with the worker. The firm discounts its future value using the interest rate adjusted for the job destruction shock $\lambda_{np}$. Since the worker may experience an extreme wealth shock with probability $\epsilon_{R}$, it changes the wage bargaining position of the worker to the firm. The future value of the firm incorporates this probability.
In the current period, $j(R,s,a)$ represents the profit earned, calculated as the output minus production costs. The worker’s productivity, $s$, is a factor in the production function $sf(k)$. The matched firm produces $sf(k)$ units of output and incurs costs for capital rental and depreciation, represented by $(r+\delta)k$, as well as wage costs, $\omega_{np}(R,s,a)$.
$J_{np}(R,s,a)$ $=\max_{k}\Big{\{}j(R,s,a)+\frac{1-\lambda^{np}}{1+r}\sum_{s^{\prime}}\pi_{ss^{\prime}}\big{[}(1-\epsilon_{R})J_{np}(R,s^{\prime},a^{\prime})+\epsilon_{R}J_{np}(R,s^{\prime},0)\big{]}\Big{\}}$ (18)
@@ 2.6 Producing p firm
Similar to a producing $np$ firm, Equation 19 describes a producing $p$ firm. The firm rents capital and discounts future profit accounting for the job destruction shock $\lambda_{p}$ and the worker’s extreme wealth shock $\epsilon_{wh}$.